# Merlin: Multi-Stage Curriculum Alignment for Multilingual Encoder-LLM Integration in Cross-Lingual Reasoning

### v3

@@ Abstract
Large language models (LLMs) excel in English but still struggle with complex reasoning in many low-resource languages (LRLs). Existing methods align LLMs with multilingual encoders, such as LangBridge and MindMerger, raising the accuracy for mid and high-resource languages, yet large performance gap remains for LRLs. We present Merlin, a model-stacking framework that iteratively refines in 2-stages based on a curriculum strategy (from general to specific where general is bilingual bitext and specific is task-specific data) and adapts only a small set of DoRA weights. On the AfriMGSM benchmark Merlin improves exact-match accuracy by +12.9 pp over MindMerger and outperforms GPT-4o-mini by 15.2 pp. It also yields consistent gains on MGSM and MSVAMP (+0.9 and +2.8 pp), demonstrating effectiveness across both low and high-resource settings.
@@ 1 Introduction
Large language models (LLMs) have achieved impressive results across a broad range of natural language processing (NLP) tasks (openai2024gpt4technicalreport; jiang2024mixtralexperts; geminiteam2024gemini15unlockingmultimodal). Despite these advances, their performance often deteriorates significantly when dealing with low-resource languages (LRLs) (adelani2024irokobenchnewbenchmarkafrican; uemura-etal-2024-afriinstruct; ojo2024goodlargelanguagemodels). This shortfall is partly due to the limited online presence of LRLs both in data availability and NLP tools, and heavy focus on high or mid-resource languages during the pretraining of these models (touvron2023llama2openfoundation; ranathunga_neural_2023; aryabumi2024aya).
To improve reasoning capabilities for LRLs, the most straightforward approach is to adapt LLMs through continual pre-training (alves2024tower; ng2025sea; buzaaba2025lugha) or post-training (ogundepo2025improving). However, these methods often depend on access to in-domain data, which is scarce for LRLs due to the lack of annotators, speakers, and linguistic experts. Additionally, fine-tuning LLMs—especially those with billions of parameters—incurs substantial computational costs and energy usage, making these approaches less feasible in resource-constrained settings (dettmers2023qloraefficientfinetuningquantized; han2024parameterefficientfinetuninglargemodels). LangBridge (yoon-etal-2024-langbridge) and MindMerger (huang2024mindmergerefficientboostingllm) address this challenge by leveraging rich multilingual representations from an external encoder as input to decoder-only LLMs. This design significantly enhances multilingual reasoning capabilities while substantially reducing both data and computational requirements. On the other hand, curriculum learning—a training paradigm in which models are exposed to data of increasing difficulty—shown to improve both convergence speed and final performance (soviany_curriculum_2022) and have been applied in recent works to enhance LLMs reasoning ability (ma-etal-2025-problem).
@@
In this paper, we introduce Multilingual Embedding-Enhanced Reasoning for Language Integration Network (Figure 1), which is abbreviated to Merlin. It is a two-stage framework that integrates insights from model stacking with question alignment. This design combines data-efficient supervised training with targeted adaptation of the LLM’s internal representations, yielding tighter encoder–decoder alignment without the cost of full-model retraining. Inspired by the principles of curriculum learning, in the first stage, we learn a lightweight mapping layer to align a strong multilingual encoder with a frozen LLM through three successive tasks starting with general domain bilingual data, followed by math domain bilingual data, and finally the task data. In the second stage, we fine-tune the LLM with parameter-efficient methods while keeping the mapping layer fixed, enabling the LLM to internalize and better exploit the learned representations during reasoning. We evaluate Merlin on various math reasoning datasets such as Multilingual Grade School Math Benchmark (MGSM) (shi2022languagemodelsmultilingualchainofthought), Multilingual Semantic VAriant Math Problems (MSVAMP) (chen2024breakinglanguagebarriersmultilingual), and AfriMGSM (adelani2024irokobenchnewbenchmarkafrican), an extension of MGSM to several low-resource African languages. Merlin surpasses the strongest baseline, MindMerger, LayAlign, SLAM and QAlign, by more than +11.0 accuracy points on AfriMGSM and delivers consistent gains on MGSM with +2.8 performance gain to MindMerger, demonstrating its effectiveness for cross-lingual mathematical reasoning in multilingual settings, including LRLs. To assess generalization beyond mathematical reasoning, we evaluate Merlin on natural language inference tasks using AfriXNLI (adelani2024irokobenchnewbenchmarkafrican), with +1.6 improvement in performance over MindMerger.
@@ 2.1 Stage I: Model Stacking
Problem Definition: Let $q$ be a target-language sentence of length $\ell$. A pretrained multilingual encoder produces $X=\operatorname{Enc}(q)\in\mathbb{R}^{\ell\times d_{\text{enc}}}$. A two-layer perceptron $f_{\sigma}:\mathbb{R}^{d_{\text{enc}}}\!\rightarrow\!\mathbb{R}^{d_{\text{llm}}}$, with parameters $\sigma$, projects these states into the LLM space, $X_{f}=f_{\sigma}(X)$.
@@
An initial bridge is established with a generic bitext from $L_{\text{tgt}}$, a generic sentence in the target language, to English (huang2024mindmergerefficientboostingllm). The decoder receives only the projected encoder states
@@
(b) Question Alignment (zhu2024questiontranslationtrainingbetter) (Align.)
@@ 2.2 Stage II: Task Specialization
Stage II internalizes the cross-lingual signal inside the decoder; enhances the integration of multilingual representations within the decoder and plays a critical role in improving performance on complex reasoning tasks. The connector $\sigma^{\ast}$, the encoder and all pretrained LLM weights remain fixed; only rank-constrained DoRA (liu2024doraweightdecomposedlowrankadaptation) matrices $\theta$ inside the decoder are activated. For every English task example $(q,a^{\star})$ the input takes the augmented form $(X_{f},T)(q)$ defined above. The adaptation parameters are obtained through
@@
Key Multilingual Approaches-Gemma2 & NLLB 600m Distilled
@@
Table 3: AfriXNLI performance on various LLMs, MindMerger, and Merlin. Metric is Accuracy. Average computed on only African languages. NLLB 600M distilled is used as the encoder.
@@ 3.1 LLMs evaluated
We evaluate two closed models(GPT-4o-mini and GPT-4o) and four open LLMs, including Aya-101 (ustun2024ayamodelinstructionfinetuned), MetaMath 7B, MetaMath-Mistral 7B (yu2023metamath), and Gemma 2 9B (gemmateam2024gemma2improvingopen). Aya-101 is only evaluated for the AfriXNLI as it is one of the strongest open model baseline for low-resource languages.
@@ 3.2 Baselines Method Compared
We evaluate our method against five strong baselines: three model merging approaches—LangBridge (yoon-etal-2024-langbridge), MindMerger (huang2024mindmergerefficientboostingllm), LayAlign (ruan2025layalignenhancingmultilingualreasoning) and two methods leveraging in-domain question translation data, QAlign-LoRA (zhu2024questiontranslationtrainingbetter) and SLAM (fan2025slamefficientmultilingualreasoning). LangBridge injects a frozen multilingual encoder before a decoder-only LLM, replacing token embeddings with linearly projected encoder states ($\approx$3M parameters) learned from 100k English instructions, and inference cost equals that of the LLM plus the multilingual encoder. QAlign-LoRA first aligns multilingual prompts via question translation and then fine-tunes lightweight LoRA adapters (<1% parameters) on English tasks. SLAM fine-tunes the six lowest feed-forward layers ($\approx$7% parameters) identified as language-sensitive on mixed English and translated data, freezing higher layers and attention weights. MindMerger connects a frozen multilingual encoder to the LLM through a two-layer MLP head ($\approx$9M parameters), trained first on translation pairs with English references, then on translated QA pairs while keeping all base weights fixed. LayAlign integrates encoder features into multiple LLM layers via a trainable layer-wise aligner, trained in two stages using translation and task data while keeping both encoder and LLM frozen.
For the encoder component in multilingual‐integration approaches, we couple both MindMerger and Merlin with the distilled 600M‐parameter NLLB encoder (nllbteam2022languageleftbehindscaling). For LangBridge, we follow its original setup: AfriTeva-v2-Large (oladipo-etal-2023-better) is used on AfriMGSM and AfriXNLI, whereas mT5-XL is employed for MGSM, because the publicly released LangBridge implementation only supports encoders that subclass the XXXEncoderModel.
@@ 3.3 Languages and Tasks
Our evaluation spans four multilingual benchmarks. AfriMGSM (adelani2024irokobenchnewbenchmarkafrican) extends MGSM with grade-school mathematics problems translated into 16 African languages, providing fine-grained coverage of more low-resource settings. AfriXNLI (adelani2024irokobenchnewbenchmarkafrican) offers natural-language inference data in the same 16 languages by translating ten domain splits from the original XNLI. The canonical MGSM corpus (shi2022languagemodelsmultilingualchainofthought) supplies math QA instances across eight mid-resource languages. Finally, MSVAMP (chen2024breakinglanguagebarriersmultilingual) introduces multilingual variants of SVAMP with 1000 algebra-word problems in eleven languages, testing cross-lingual reasoning beyond GSM-style arithmetic.
@@ 3.4 Experimental Settings
Sub-stage Ia (Mapping). For every language we sample 9 000 general-domain English–target sentence pairs from the NLLB training corpus (nllbteam2022languageleftbehindscaling). These bitext pairs supply a coarse bilingual anchor for the connector.
@@
To align non-English prompts with English semantics, we translate 20 000 English questions with NLLB 200-3.3B and sample 3 000 translated pairs per language. For the mathematics track (AfriMGSM and MGSM) the source questions come from MetaMathQA (yu2024metamathbootstrapmathematicalquestions). For the NLI track (AfriXNLI) we instead sample premise–hypothesis pairs from translated MultiNLI (williams2018broadcoveragechallengecorpussentence), concatenate them as the “question”, translate, and again retain 3 000 pairs.
@@
We fine-tune the LLM (via PEFT) on a second sample of 3 000 training instances per language drawn from the same translated sources as above. Using a separate set of examples helps reduce the risk of overlap with Sub-stage Ic and ensures there is no data leakage into the evaluation sets. [^1] The detailed training configuration is stated on Appendix
@@ 4 Experimental Results
We assessed Merlin on both math reasoning and NLI benchmarks, comparing it against open LLMs, recent multilingual integration methods, and the strongest available closed models. For math reasoning, we use two datasets: MSGM and MSVAMP. To assess performance on low-resource African languages across both tasks, we employ AfriMGSM and AfriXNLI. All results are reported as exact-match or accuracy; higher is better.
@@
Table 1 presents results on MGSM and its algebraic variant MSVAMP. Using the identical Gemma-2 9B backbone, Merlin achieves $76.2\,\%$ average accuracy on MGSM, exceeding MindMerger by +2.8+2.8 points and SLAM by +22.8+22.8 points. [^2] LangBridge and QAlign scores are omitted in Table 1 because their publicly released implementations do not support the full language set. Improvements are consistent across both mid-resource languages (Bn/Th/Sw, +13.7+13.7 over Gemma-2) and high-resource languages (+7.7+7.7 over MetaMath-Mistral). On MSVAMP, Merlin retains its lead with $79.2\,\%$ average accuracy, outperforming MindMerger by +0.9+0.9 points and exceeding the best open baseline by more than $15$ points. These gains indicate that the two-stage alignment not only transfers arithmetic skills but also generalises to algebraic word problems.
@@
Table 2 focuses on sixteen truly low-resource African languages. With the strong Gemma-2 9B+NLLB-600M configuration, Merlin attains $50.6\,\%$ average accuracy, surpassing MindMerger by +12.9+12.9 points, SLAM by +11.6+11.6 points, and QAlign-LoRA by +16.9+16.9 points. Notably, Merlin outruns the closed GPT-4o-mini by +15.2+15.2 points and comes within $2.0$ points of GPT-4o despite the latter’s much larger parameter budget. Ablation confirms (Table 4) that the task-aware augmentation sub-stage is decisive, yielding gains of more than ten points when added to the connector training.
Reasoning capabilities generalizes to NLU tasks
Table 3 reports accuracy on sixteen African languages plus English and French on NLI task. Merlin attains $68.7\,\%$ average accuracy on the African subset, improving on MindMerger by +1.6+1.6 points and on the base Gemma–2 9B model by +28.4+28.4 points, and marginally surpassing the closed GPT‑4o ($64.3\,\%$). The gains are most pronounced for Amharic (+5.6), Oromo (+4.6), Hausa (+2.5), and Kinyarwanda (+2.0), confirming the value of the cross‑lingual mapping curriculum in truly low‑resource settings. The overall improvement across thirteen of the sixteen African languages indicates that Merlin successfully redistributes modelling capacity toward languages that benefit the most from additional cross‑lingual supervision, while maintaining competitive performance on those already well supported.

### v4

@@ Abstract
Large language models (LLMs) excel in English but still struggle with complex reasoning in many low-resource languages (LRLs). Existing methods align LLMs with multilingual encoders, such as LangBridge and MindMerger, raising the accuracy for mid and high-resource languages, yet large performance gap remains for LRLs. We present Merlin, a model-stacking framework that iteratively refines in two-stages based on a curriculum strategy (from general to specific where general is bilingual bitext and specific is task-specific data) and adapts only a small set of DoRA weights. On the AfriMGSM benchmark Merlin improves exact-match accuracy by +12.9 pp over MindMerger and outperforms GPT-4o-mini by 15.2 pp. It also yields consistent gains on MGSM and MSVAMP (+0.9 and +2.8 pp), demonstrating effectiveness across both low and high-resource settings.  [^1] The implementation of Merlin is publicly available at https://github.com/LLMforLRL/MERLIN.
@@ 1 Introduction
Large language models (LLMs) have achieved impressive results across a broad range of natural language processing (NLP) tasks (OpenAI et al., 2024; Jiang et al., 2024; Team et al., 2024a). Despite these advances, their performance often deteriorates significantly when dealing with low-resource languages (LRLs) (Adelani et al., 2024b; Uemura et al., 2024; Ojo et al., 2024). This shortfall is partly due to the limited online presence of LRLs both in data availability and NLP tools, and heavy focus on high or mid-resource languages during the pretraining of these models (Touvron et al., 2023; Ranathunga et al., 2023; Aryabumi et al., 2024).
To improve reasoning capabilities for LRLs, the most straightforward approach is to adapt LLMs through continual pre-training (Alves et al., 2024; Ng et al., 2025; Buzaaba et al., 2025) or post-training (Ogundepo et al., 2025). However, these methods often depend on access to in-domain data, which is scarce for LRLs due to the lack of annotators, speakers, and linguistic experts. Additionally, fine-tuning LLMs—especially those with billions of parameters—incurs substantial computational costs and energy usage, making these approaches less feasible in resource-constrained settings (Dettmers et al., 2023; Han et al., 2024). LangBridge (Yoon et al., 2024) and MindMerger (Huang et al., 2024) address this challenge by leveraging rich multilingual representations from an external encoder as input to decoder-only LLMs. This design significantly enhances multilingual reasoning capabilities while substantially reducing both data and computational requirements. On the other hand, curriculum learning—a training paradigm in which models are exposed to data of increasing difficulty—has been shown to improve both convergence speed and final performance (Soviany et al., 2022) and has been applied in recent works to enhance LLMs reasoning ability (Ma et al., 2025).
@@
In this paper, we introduce Multilingual Embedding-Enhanced Reasoning for Language Integration Network (Figure 1), which is abbreviated to Merlin. It is a two-stage framework that integrates insights from model stacking with question alignment. This design combines data-efficient supervised training with targeted adaptation of the LLM’s internal representations, yielding tighter encoder–decoder alignment without the cost of full-model retraining. Inspired by the principles of curriculum learning, in the first stage, we learn a lightweight mapping layer to align a strong multilingual encoder with a frozen LLM through three successive tasks starting with general domain bilingual data, followed by math domain bilingual data, and finally the task data. In the second stage, we fine-tune the LLM with parameter-efficient methods while keeping the mapping layer fixed, enabling the LLM to internalize and better exploit the learned representations during reasoning. We evaluate Merlin on various math reasoning datasets such as Multilingual Grade School Math Benchmark (MGSM) (Shi et al., 2022), Multilingual Semantic VAriant Math Problems (MSVAMP) (Chen et al., 2024), and AfriMGSM (Adelani et al., 2024b), an extension of MGSM to several low-resource African languages. Merlin surpasses the strongest baseline, MindMerger, LayAlign, SLAM and QAlign, by more than +11.0 accuracy points on AfriMGSM and delivers consistent gains on MGSM with +2.8 performance gain to MindMerger, demonstrating its effectiveness for cross-lingual mathematical reasoning in multilingual settings, including LRLs. To assess generalization beyond mathematical reasoning, we evaluate Merlin on natural language inference tasks using AfriXNLI (Adelani et al., 2024b), with +1.6 improvement in performance over MindMerger.
@@ 2.1 Stage I: Model Stacking
Problem Definition:
Let $q$ be a target-language sentence of length $\ell$. A pretrained multilingual encoder produces $X=\operatorname{Enc}(q)\in\mathbb{R}^{\ell\times d_{\text{enc}}}$. A two-layer perceptron $f_{\sigma}:\mathbb{R}^{d_{\text{enc}}}\!\rightarrow\!\mathbb{R}^{d_{\text{llm}}}$, with parameters $\sigma$, projects these states into the LLM space, $X_{f}=f_{\sigma}(X)$.
@@
An initial bridge is established with a generic bitext from $L_{\text{tgt}}$, a generic sentence in the target language, to English (Huang et al., 2024). The decoder receives only the projected encoder states
@@
(b) Question Alignment (Zhu et al., 2024) (Align.)
@@ 2.2 Stage II: Task Specialization
Stage II internalizes the cross-lingual signal inside the decoder; enhances the integration of multilingual representations within the decoder and plays a critical role in improving performance on complex reasoning tasks. The connector $\sigma^{\ast}$, the encoder and all pretrained LLM weights remain fixed; only rank-constrained DoRA (Liu et al., 2024) matrices $\theta$ inside the decoder are activated. For every English task example $(q,a^{\star})$ the input takes the augmented form $(X_{f},T)(q)$ defined above. The adaptation parameters are obtained through
@@
2.3 Multi-stage instead of End-to-end design justification.
Our multi-stage design is motivated by stability, data efficiency, and computational practicality. End-to-end training of the encoder $\to$ connector $\to$ decoder stack using limited LRL supervision tends to overfit and destabilize the coarse bilingual alignment learned from general bitext. Moreover, Stage Ia is task-agnostic and can be reused across math and NLI, whereas StageIb, Ic, and Stage II are lightweight task-specific steps; joint end-to-end training would entangle these and make reuse difficult.
@@
Key Multilingual Approaches-Gemma2 & NLLB-600m Distilled
@@
Table 3: AfriXNLI performance on various LLMs, MindMerger, and Merlin. Metric is Accuracy. Average computed on only African languages. NLLB-600M distilled is used as the encoder.
@@ 3.1 LLMs evaluated
We evaluate two closed models(GPT-4o-mini and GPT-4o) and four open LLMs, including Aya-101 (Üstün et al., 2024), MetaMath 7B, MetaMath-Mistral 7B (Yu et al., 2023), and Gemma 2 9B (Team et al., 2024b). Aya-101 is only evaluated for the AfriXNLI as it is one of the strongest open model baselines for low-resource languages.
Our choice of MetaMath 7B/Mistral and Gemma 2 9B is based on their strong English mathematical reasoning abilities, which is crucial because Merlin’s goal is to transfer existing reasoning competence from English to target languages. Many general-purpose models (e.g., LLaMA 3.1 Grattafiori et al. (2024)) underperform on GSM-style math Adelani et al. (2024b), leaving little reasoning skill to transfer. Overall, Merlin is backbone-agnostic and can be applied to other LLMs in principle.
@@ 3.2 Baselines Method Compared
We evaluate our method against five strong baselines: three model merging approaches—LangBridge (Yoon et al., 2024), MindMerger (Huang et al., 2024), LayAlign (Ruan et al., 2025) and two methods leveraging in-domain question translation data, QAlign-LoRA (Zhu et al., 2024) and SLAM (Fan et al., 2025). LangBridge injects a frozen multilingual encoder before a decoder-only LLM, replacing token embeddings with linearly projected encoder states ($\approx$3M parameters) learned from 100k English instructions, and inference cost equals that of the LLM plus the multilingual encoder. QAlign-LoRA first aligns multilingual prompts via question translation and then fine-tunes lightweight LoRA adapters (<1% parameters) on English tasks. SLAM fine-tunes the six lowest feed-forward layers ($\approx$7% parameters) identified as language-sensitive on mixed English and translated data, freezing higher layers and attention weights. MindMerger connects a frozen multilingual encoder to the LLM through a two-layer MLP head ($\approx$9M parameters), trained first on translation pairs with English references, then on translated QA pairs while keeping all base weights fixed. LayAlign integrates encoder features into multiple LLM layers via a trainable layer-wise aligner, trained in two stages using translation and task data while keeping both encoder and LLM frozen.
For the encoder component in multilingual‐integration approaches, we couple both MindMerger and Merlin with the distilled 600M‐parameter NLLB encoder (Team et al., 2022). For LangBridge, we follow its original setup: AfriTeva-v2-Large (Oladipo et al., 2023) is used on AfriMGSM and AfriXNLI, whereas mT5-XL is employed for MGSM, since the publicly released LangBridge implementation only supports encoders of the class XXXEncoderModel (e.g., MT5EncoderModel).
@@ 3.3 Languages and Tasks
Our evaluation spans four multilingual benchmarks. AfriMGSM (Adelani et al., 2024b) extends MGSM with grade-school mathematics problems translated into 16 African languages, providing fine-grained coverage of more low-resource settings. AfriXNLI (Adelani et al., 2024b) offers natural-language inference data in the same 16 languages by translating ten domain splits from the original XNLI. The canonical MGSM corpus (Shi et al., 2022) supplies math QA instances across eight mid-resource languages. Finally, MSVAMP (Chen et al., 2024) introduces multilingual variants of SVAMP with 1000 algebra-word problems in eleven languages, testing cross-lingual reasoning beyond GSM-style arithmetic.
@@ 3.4 Experimental Settings
Sub-stage Ia (Mapping). For every language we sample 9 000 general-domain English–target sentence pairs from the NLLB training corpus (Team et al., 2022). These bitext pairs supply a coarse bilingual anchor for the connector.
@@
To align non-English prompts with English semantics, we translate 20 000 English questions with NLLB 200-3.3B and sample 3 000 translated pairs per language. For the mathematics track (AfriMGSM and MGSM) the source questions come from MetaMathQA (Yu et al., 2024). For the NLI track (AfriXNLI) we instead sample premise–hypothesis pairs from translated MultiNLI (Williams et al., 2018), concatenate them as the “question”, translate, and again retain 3 000 pairs.
@@
We fine-tune the LLM (via PEFT) on a second sample of 3 000 training instances per language drawn from the same translated sources as above. Using a separate set of examples helps reduce the risk of overlap with Sub-stage Ic and ensures there is no data leakage into the evaluation sets. [^2] The detailed training configuration is provided on Appendix
Alignment of training data across methods.
Merlin and all baselines are trained on the same underlying supervision. For math tasks, all methods (MindMerger, QAlign, SLAM, Merlin) use the same question-level translations derived from MetaMathQA; for AfriXNLI, all methods use premise–hypothesis pairs derived from MultiNLI. The only difference lies in how each method utilizes this data (e.g., which module is trained at which stage).
@@ 4 Experimental Results
We assessed Merlin on both math reasoning and NLI benchmarks, comparing it against open LLMs, recent multilingual integration methods, and the strongest available closed models. For math reasoning, we use two datasets: MGSM and MSVAMP. To assess performance on low-resource African languages across both tasks, we employ AfriMGSM and AfriXNLI. All results are reported as exact-match or accuracy; higher is better.
@@
Table 1 presents results on MGSM and its algebraic variant MSVAMP. Using the identical Gemma 2 9B backbone, Merlin achieves $76.2\,\%$ average accuracy on MGSM, exceeding MindMerger by +2.8+2.8 points and SLAM by +22.8+22.8 points. [^3] LangBridge and QAlign scores are omitted in Table 1 because their publicly released implementations do not support the full language set. Improvements are consistent across both mid-resource languages (Bn/Th/Sw, +13.7+13.7 over Gemma 2) and high-resource languages (+7.7+7.7 over MetaMath-Mistral). On MSVAMP, Merlin retains its lead with $79.2\,\%$ average accuracy, outperforming MindMerger by +0.9+0.9 points and exceeding the best open baseline by more than $15$ points. These gains indicate that the two-stage alignment not only transfers arithmetic skills but also generalises to algebraic word problems.
@@
Table 2 focuses on sixteen truly low-resource African languages. With the strong Gemma 2 9B+NLLB-600M configuration, Merlin attains $50.6\,\%$ average accuracy, surpassing MindMerger by +12.9+12.9 points, SLAM by +11.6+11.6 points, and QAlign-LoRA by +16.9+16.9 points. Notably, Merlin outruns the closed GPT-4o-mini by +15.2+15.2 points and comes within $2.0$ points of GPT-4o despite the latter’s much larger parameter budget. Ablation confirms (Table 4) that the task-aware augmentation sub-stage is decisive, yielding gains of more than ten points when added to the connector training.
Reasoning capabilities generalize to NLU tasks
Table 3 reports accuracy on sixteen African languages plus English and French on NLI task. Merlin attains $68.7\,\%$ average accuracy on the African subset, improving on MindMerger by +1.6+1.6 points and on the base Gemma 2 9B model by +28.4+28.4 points, and marginally surpassing the closed GPT‑4o ($64.3\,\%$). The gains are most pronounced for Amharic (+5.6), Oromo (+4.6), Hausa (+2.5), and Kinyarwanda (+2.0), confirming the value of the cross‑lingual mapping curriculum in truly low‑resource settings. The overall improvement across thirteen of the sixteen African languages indicates that Merlin successfully redistributes modelling capacity toward languages that benefit the most from additional cross‑lingual supervision, while maintaining competitive performance on those already well supported.